OpenStudy (mathmath333):

Harsh makes a fixed deposit of 20000 with the bank of india for a period of 3 years , if the rate of interest is 13 % simple interest charged half yearly, what will the amount he get after 42 months ?

OpenStudy (anonymous):

Take the interest rate from the 20000 7 times.

OpenStudy (anonymous):

reminds me of leonard Cohen..everybody knows!

OpenStudy (anonymous):

Because, if it is charged half-yearly, it will be charged every six months, because there are 12 months in a year. It wants the total charge after 42 months. 42/7=6. The rate of interest is 13%. Multiply what you get after taking from the 20000 once with the rate of interest, & multiply that by 7.

OpenStudy (mathmath333):

@ sith ans giggle it is mentioned simple interst not compond interset in question

OpenStudy (anonymous):

Does that make sense @mathmath333

OpenStudy (anonymous):

Principal sum = 20,000

OpenStudy (anonymous):

n = 3

OpenStudy (anonymous):

r = 13%

OpenStudy (mathmath333):

@Brostep0s continue pleease

OpenStudy (anonymous):

Do you know how to apply the rate of interest to the principal sum?

OpenStudy (mathmath333):

i=prt

OpenStudy (mathmath333):

please continue @Brostep0s

OpenStudy (anonymous):

Apply that one time. Please show your answer.

OpenStudy (mathmath333):

u mean i=20000*0.13*3 ?

OpenStudy (anonymous):

*7, at the end, but yes.

OpenStudy (mathmath333):

why 7?

OpenStudy (mathmath333):

is final answer \(54600\) @Brostep0s

OpenStudy (anonymous):

can you please show calculation @mathmath333

OpenStudy (mathmath333):

http://www.wolframalpha.com/input/?i=%3D20000*0.13*3+*7

OpenStudy (anonymous):

This is incorrect. The correct answer is 18200 charged over 42 months. You made your error when you multiplied 3 into the equation.

OpenStudy (anonymous):

@Brostep0s you cant say that until you show calculation

OpenStudy (anonymous):

? i=20000*.13*7= 18200 total charge. {You made your error when you multiplied 3 into the equation} i=20000*.13{*3}*7.

OpenStudy (mathmath333):

yes the \(18200\) in not given correct in book

OpenStudy (anonymous):

What are your choices, & I'll see what i did wrong.

OpenStudy (anonymous):

Keep the three included. i missed that piece of information.

OpenStudy (mathmath333):

a. 27800 b. 28100 c. 29100 d. 28500

OpenStudy (anonymous):

@triciaal Help me with what I did wrong, if you will.

OpenStudy (triciaal):

Simple Interest = Principal in $ * rate as % * time in years I = P R T/100

OpenStudy (anonymous):

nice to see you @triciaal